OpenStudy (anonymous):

What is the solution set of |x – 6| = 4? 10 10 or –10 –2 or –10 2 or 10

OpenStudy (helder_edwin):

use this theorem |a|=b if and only if a=b or a=b

OpenStudy (helder_edwin):

sorry the socond is a=-b

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

2 or 10

OpenStudy (helder_edwin):

solve x-6=4 or x-6=-4

OpenStudy (helder_edwin):

yes that is the solution

OpenStudy (anonymous):

2 or 10

OpenStudy (anonymous):

x-6=4 or x-6=-4 x=10 or x=2